# An open-source heuristic to reboot 2D nesting research

### v2

@@ 9.1 Exploration phase
The separation procedure (Algorithm 9) attempts to convert this infeasible solution into a feasible one. If successful, the returned solution $s$ replaces the incumbent solution $s^{*}$, after which the strip is once again shrunk. If unsuccessful, $s$ represents an infeasible local optimum that is added to a pool of solutions $\mathcal{S}$. For the next separation attempt, a random solution $\hat{s}$ is chosen from $\mathcal{S}$, with those closer to feasibility (Equation 8) having a greater chance of being selected. A new starting solution $s^{\prime}$ is then created by swapping two large items from $\hat{s}$. This ensures $s^{\prime}$ inherits many high-quality placements from $\hat{s}$ while, at the same time, being sufficiently disrupted so it can escape from the local optimum. This procedure continues until time limit $\texttt{TL}_{x}$ is reached, at which point the exploration phase returns the best feasible solution found $s^{*}$.
@@ 9.2 Compression phase
3:   $r\leftarrow\texttt{R}_{c}^{s}+(\texttt{R}_{c}^{e}-\texttt{R}_{c}^{s})\cdot(\tau/\texttt{TL}_{c})$
@@ 10 Implementation
The implementation of our 2D nesting algorithm is written in Rust and called sparrow. It is publicly available at: https://github.com/JeroenGar/sparrow. The codebase closely follows the structure of this paper, making it straightforward to understand how sparrow relates to the algorithms presented here. Throughout this project, we have strived for maximum computational performance, robustness and leanness of both the algorithm and its implementation.
@@
So far, this limitation concerning exact fits only surfaces in artificial academic instances featuring simple shapes and vertices with integer coordinates. Conditions that, generally, do not occur in real-world applications.
@@ 11.1 Experimental setup
We used a combination of manual tuning and hyperparameter optimization (bergstra2013making) to identify a suitable set of parameters for the algorithm. These parameters were optimized for the maximum expected performance across all instances from the academic benchmark set when subject to a 20-minute time limit. This single set of parameters was used for every instance in our experiments.
@@ 11.2 Performance analysis
The expected quality $\mathbb{E}(\rho)$ and interquartile range of sparrow’s solutions in comparison to the academic SotA algorithms for the 2DISPP and one open-source implementation are presented in Table 2. The distributions of the solutions from Table 2 are visualized as violin plots in Figure 9. One important thing to note is that the width of the strip for some of the most simple instances – SHAPES0, SHAPES1, SHAPES2, FU, JAKOBS1 and JAKOBS2 – was inflated by 0.01%. This minuscule inflation should not affect which configurations of items are possible, but suffices to avoid the issue concerning exact fits (described in Section 10), which would otherwise occur in these artificial instances. For academic purposes, sparrow’s best solution ever produced for each instance, alongside the previous best-known solution, is presented in Table 3 and shown in Figure 10. It is important to note that these ‘best’ solutions, although conceived under the same experimental conditions and time limit, are not necessarily elements of the sets of 100 independent runs from which the expected densities in Table 2 have been derived.
@@
Table 2: The expected solution quality $\mathbb{E}(\rho)$ on academic benchmarks obtained by sparrow, four algorithms from the literature: ROMA - sato2019raster, GCS - elkeran2013new, FLD - wang2017flexible, ELS - leung2012extended and one open-source implementation: PS - fontan. For sparrow, quartiles (Q1, Q3) and the margin of error ($\pm$) on $\mathbb{E}$, defined as the half-width of the 95% confidence interval computed via non-parametric bootstrapping, are reported. The reported values for sparrow result from 100 independent runs, each with a time limit of 20 minutes. PS is a deterministic algorithm and was run once under the same conditions as sparrow. For each instance, the best $\mathbb{E}(\rho)$ is highlighted in bold.
@@
The algorithms composing the academic SotA report only two values per instance: the best and the expected density from a limited number of runs. This limits our ability to provide detailed statistical or visual performance comparisons against them. However, to support future in-depth comparisons against sparrow, we provide all solution data and a complete reproducibility guide in the repository.
@@ 12 Real-world benchmarks & future research directions
In Section 11.2, we demonstrated that sparrow is capable of consistently producing high-quality solutions across the entire set of traditional academic benchmark instances. This set, however, is dated and contains many instances that were constructed artificially, featuring simple polygonal shapes with vertices located at integer coordinates. For SWIM, TROUSERS and SHIRTS, arguably the instances that most reflect real-world conditions, sparrow makes a significant leap forward. However, these instances are few in number and only represent a single application domain: the garment industry. We therefore take the opportunity to extend the set of academic benchmarks to better capture the breadth and intricacies of real-world applications.
@@ 13 Conclusions and reflections
This paper introduced a heuristic for the 2D irregular strip packing problem (2DISPP) that redefines both the academic and open-source state of the art. Our implementation, sparrow, is publicly available at: https://github.com/JeroenGar/sparrow. The core idea of the algorithm is to decompose the 2DISPP into a sequence of feasibility problems, where the strip is converted into a container with fixed dimensions and the items are temporarily allowed to collide with each other. A local-search algorithm then iteratively moves colliding items around in an attempt to gradually resolve them all and eventually reach feasibility. Although we focus on the 2DISPP, most of the heuristic’s components operate on the feasibility problem and can easily be incorporated into solution strategies for other 2D nesting problems such as bin packing or knapsack problems.
The main contributions of this work are fourfold: (i) a reproducible and substantial improvement in solution quality across both academic and open-source contexts, (ii) a highly-optimized and robust implementation ready for deployment in practice or as a base for future research, (iii) a number of foundational algorithmic ideas that can be iteratively improved upon or adapted to other problem variants and (iv) a new set of academic benchmark instances designed to properly reflect real-world nesting applications and uncover weaknesses in our approach.
To be frank, the scale of improvements with respect to the previous SotA demonstrated in the computational experiments (Section 11) should not have been possible on such a well-studied problem using benchmark instances that are over 25 years old. On the most challenging instance, SWIM, sparrow’s expected and best solution quality are both $\sim$5% more compact than the previous best results. Such a margin of improvement confirms that stagnation with respect to 2DISPP research was not a result of it already being solved to near optimality.
@@
Reproducibility is a cornerstone of the scientific method. Without it, researchers are left in an unhealthy and inefficient environment: forced to reinvent the wheel without the means to increment on prior work or even validate its results. We strongly believe that an academic paper coupled with a high-quality and open-source implementation delivers value significantly exceeding the sum of its parts. Thankfully, the ethos we are advocating is not swimming against the current. The field of Operations Research is undergoing a gradual cultural shift toward open-source practices (petropoulos2024operational). vidal2022hybrid, for example, recently published an open-source implementation of their state-of-the-art algorithm for the capacitated vehicle routing problem. Many of the concerns we raise here are in fact echoing sentiments already expressed in vidal2022hybrid.

### v3

@@ 9.1 Exploration phase
The separation procedure (Algorithm 9) attempts to convert this infeasible solution into a feasible one. If successful, the returned solution $s$ replaces the incumbent solution $s^{*}$, after which the strip is once again shrunk. If unsuccessful, $s$ represents an infeasible local optimum that is added to a pool of solutions $\mathcal{S}$. For the next separation attempt, a random solution $\hat{s}$ is chosen from $\mathcal{S}$, with those closer to feasibility (Expression 8) having a greater chance of being selected. A new starting solution $s^{\prime}$ is then created by swapping two large items from $\hat{s}$. This ensures $s^{\prime}$ inherits many high-quality placements from $\hat{s}$ while, at the same time, being sufficiently disrupted so it can escape from the local optimum. This procedure continues until time limit $\texttt{TL}_{x}$ is reached, at which point the exploration phase returns the best feasible solution found $s^{*}$.
@@ 9.2 Compression phase
3:   $r\leftarrow\texttt{R}_{c}^{s}+(\texttt{R}_{c}^{e}-\texttt{R}_{c}^{s})\cdot(\tau/\texttt{TL}_{c})$ $\triangleright$ Table 1
@@ 10 Implementation
Our implementation of the proposed 2D nesting algorithm is written in Rust and called sparrow. It is publicly available at: https://github.com/JeroenGar/sparrow. The codebase closely follows the structure of this paper, with all introduced algorithms being referenced directly in the source code. This makes it straightforward to understand how exactly sparrow maps to the algorithms presented here. Throughout this project, we have strived for maximum computational performance, robustness and leanness of both the algorithm and its implementation.
@@
At present, this limitation concerning exact fits only surfaces in artificial academic instances featuring simple shapes and vertices with integer coordinates. In fact, real-world applications often require a minimum separation distance between items to account for manufacturing tolerances or cutting tool widths. This is a feature that jagua-rs already supports natively.
@@ 11.1 Experimental setup
We used a combination of manual tuning and hyperparameter optimization [bergstra2013making] to identify a suitable set of parameters for the algorithm. These parameters were optimized for the maximum expected performance across all instances from the academic benchmark set when subject to a 20-minute time limit. This single set of parameters was used for every instance in our experiments.
@@ 11.2 Performance analysis
The expected quality $\mathbb{E}(\rho)$ and interquartile range of sparrow’s solutions in comparison to the academic state of the art algorithms for the 2DISPP and one open-source implementation are presented in Table 2. The distributions of the solutions from Table 2 are visualized as violin plots in Figure 9. We performed 100 independent runs for each instance to ensure our reported expected values are statistically robust, with sufficiently narrow confidence intervals. It is important to note that the width of the strip for some of the most simple instances – SHAPES0, SHAPES1, SHAPES2, FU, JAKOBS1 and JAKOBS2 – was inflated by 0.01%. This minuscule inflation does not affect the possible feasible item configurations, but suffices to avoid the issue concerning exact fits (described in Section 10), which would otherwise occur in these artificial instances. For academic purposes, sparrow’s best solution ever produced for each instance, alongside the previous best-known solution, is presented in Table 3 and shown in Figure 10. It is important to note that these ‘best’ solutions, although conceived under the same experimental conditions and time limit, are not necessarily elements of the sets of 100 independent runs from which the expected densities in Table 2 have been derived.
@@
Table 2: The expected solution quality $\mathbb{E}(\rho)$ on academic benchmarks obtained by sparrow, four algorithms from the literature: ROMA - sato2019raster, GCS - elkeran2013new, FLD - wang2017flexible, ELS - leung2012extended and one open-source implementation: PS - fontan. For sparrow, quartiles (Q1, Q3) and the margin of error ($\pm$) on $\mathbb{E}$, defined as the half-width of the 95% confidence interval computed via non-parametric bootstrapping, are reported. The reported values for sparrow result from 100 independent runs, each with a time limit of 20 minutes. Values for ROMA, GCS, FLD and ELS were taken from their respective publications. PS is a deterministic algorithm and was run once under the same conditions as sparrow. For each instance, the best $\mathbb{E}(\rho)$ is highlighted in bold.
@@
The algorithms composing the academic state of the art report only two values per instance: the best and the expected density from a limited number of runs. This limits our ability to provide detailed statistical or visual performance comparisons against them. However, to support future in-depth comparisons against sparrow, we provide all solution data and a complete reproducibility guide in the repository.
@@ 12 Real-world benchmarks & future research directions
Section 11.2 demonstrated that sparrow is capable of consistently producing high-quality solutions across the entire set of traditional academic benchmark instances. This set, however, is dated and contains many instances that were constructed artificially, featuring simple polygonal shapes with vertices located at integer coordinates. For SWIM, TROUSERS and SHIRTS, arguably the instances that most reflect real-world conditions, sparrow makes a significant leap forward. However, these instances are few in number and only represent a single application domain: the garment industry. We therefore take the opportunity to extend the set of academic benchmark instances to better capture the breadth and intricacies of real-world applications.
@@ 13 Conclusions and reflections
This paper introduced a heuristic for the 2D irregular strip packing problem (2DISPP) that redefines both the academic and open-source state of the art. Our implementation, sparrow, is publicly available at: https://github.com/JeroenGar/sparrow. The core idea of the algorithm is to decompose the 2DISPP into a sequence of feasibility problems, where the strip is converted into a container with fixed dimensions and the items are temporarily allowed to collide with each other. A local-search algorithm iteratively moves colliding items around in an attempt to gradually resolve them all and eventually reach feasibility. Although we focus on the 2DISPP, most of the heuristic’s components operate on the feasibility problem and can thus easily be incorporated into solution strategies for other 2D nesting problems such as bin packing or knapsack problems.
The main contributions of this work are fourfold: (i) a reproducible and substantial improvement in solution quality across both academic and open-source contexts, (ii) a highly-optimized and robust implementation ready for deployment in practice or as a base for future research, (iii) a number of foundational algorithmic ideas that can be iteratively improved upon or adapted to other problem variants and (iv) a new set of academic benchmark instances designed to properly reflect real-world nesting applications and uncover weaknesses of the approach.
To be frank, the scale of improvements with respect to the previous state of the art demonstrated in the computational experiments (Section 11) should not have been possible on such a well-studied problem using benchmark instances that are over 25 years old. For the most challenging instance, SWIM, sparrow’s expected and best solution are both $\sim$5% more compact than the previous best results. Such a margin of improvement confirms that stagnation with respect to 2DISPP research was not a result of it already being solved to near optimality.
@@
Reproducibility is a cornerstone of the scientific method. Without it, researchers are left in an unhealthy and inefficient environment: forced to reinvent the wheel without the means to increment on prior work or even validate its results. We strongly believe that an academic paper coupled with a high-quality and open-source implementation delivers value significantly exceeding the sum of its parts. Thankfully, the ethos we are advocating is not swimming against the current. The field of Operations Research is undergoing a gradual cultural shift toward open-source practices [petropoulos2024operational]. vidal2022hybrid, for example, recently published an open-source implementation of their state-of-the-art algorithm for the capacitated vehicle routing problem. Many of the concerns we raise here are in fact echoing sentiments already expressed by vidal2022hybrid.